An object is projected upward from the top of a tower. Its distance in feet above the ground after t seconds is given by s(t)=-16t+64t+80. How many seconds will it take to reach ground level?
Answer by stanbon(75887)

An object is projected upward from the top of a tower. Its distance in feet above the ground after t seconds is given by s(t)=-16t^2+64t+80. How many seconds will it take to reach ground level?
-------------------
When it reaches the ground s(t)=zero.
Solve -16t^2+64t+80=0
-16(t^2-4t-5)=0
(t-5)(t+1)=0
t=5 seconds or t=-1 second
Only t=5 seconds makes any sense for this problem.
=============
Cheers,
Stan H.